OpenStudy (anonymous):

evaluate integral of (2-cosx + sinx)/(sin^2x) dx

OpenStudy (anonymous):

\[\int\limits \frac{2-\cos(x)+\sin(x)}{\sin^2(x)}dx=2\int\limits \csc^2(x)dx-\int\limits \cot(x)\csc(x)dx+\int\limits \csc(x)dx\] \[-2\cot(x)+\csc(x)-\ln|\csc(x)+\cot(x)|+C\]

OpenStudy (anonymous):

where did the 2integral csc^2 (x) dx come from? Its correct by the way...

OpenStudy (anonymous):

Because you have: \[2 \int\limits \frac{dx}{\sin^2(x)}\] And we know: \[\frac{1}{\sin(x)}=\csc(x)\]

OpenStudy (anonymous):

you saved me

OpenStudy (anonymous):

Haha, you're welcome :P

OpenStudy (anonymous):

quick question, where did you get \[\int\limits_{}^{}\cot(x)\csc(x)dx\] from?

OpenStudy (anonymous):

\[\cot(x)=\frac{\cos(x)}{\sin(x)}\]